1.5 Solve the equation below using Inverse method (10 marks)

"\begin{bmatrix}
 1 & 3 & 0 \\
 0 & 0.5 & 1\\
0.05 & 0 & 1\\

 \end{bmatrix}" "\begin{bmatrix}
 x \\
 y\\
z\\
\end{bmatrix}" ="\begin{bmatrix}
 4 \\
 1\\
3\\
\end{bmatrix}"


1
Expert's answer
2022-04-16T04:14:05-0400

"A=\begin{bmatrix}
 1 & 3 & 0\\
 0 & 0.5 & 1\\
0.05 & 0 & 1
\end{bmatrix},X=\begin{bmatrix}
 x \\
 y \\
z
\end{bmatrix},
B=\begin{bmatrix}
 4 \\
 1\\
3
\end{bmatrix}."

Matrix 𝑋 satisfying the equation 𝐴𝑋=𝐵

 is given by "X=A^{-1} B."

Let us begin by finding the inverse of the 3×3 matrix .

"\text{det} A=a_{11}|A_{11}|-a_{12}|A_{12}|+a_{13}|A_{13}|="

"=1\cdot\begin{vmatrix} 0.5& 1\\0&1\end{vmatrix}-3\cdot\begin{vmatrix} 0& 1\\0.05&1\end{vmatrix}+0\cdot\begin{vmatrix} 0& 0.5\\0.05&0\end{vmatrix}="

"=1\cdot(0.5\cdot1-1\cdot0)-3\cdot(0\cdot1-1\cdot0.05)+0=0.65;"


"+|A_{11 } |=+\begin{vmatrix}0.5 & 1 \\0 & 1\end{vmatrix}=0.5\cdot1-1\cdot0=0.5;"


"-|A_{12} |=-\begin{vmatrix}0 & 1 \\0.05& 1\end{vmatrix}=-(0\cdot1-1\cdot0.05)=0.05;"


"+|A_{13} |=+\begin{vmatrix}0 & 0.5\\0.05 & 0\end{vmatrix}=-0.025;\\ \ \\
-|A_{21} |=-\begin{vmatrix}3 & 0 \\0& 1\end{vmatrix}=-3;\\"

"+|A_{22} |=+\begin{vmatrix}1& 0\\0.05 & 1\end{vmatrix}=1;\\ \ \\
-|A_{23} |=-\begin{vmatrix}1 & 3 \\0.05& 0\end{vmatrix}=0.15;"


"+|A_{31} |=+\begin{vmatrix}3 & 0\\0 & 1\end{vmatrix}=3;\\ \ \\
-|A_{32} |=-\begin{vmatrix}1 & 0 \\0& 1\end{vmatrix}=-1;\\ \ \\
+|A_{33} |=+\begin{vmatrix}1 & 3\\0 & 0.5\end{vmatrix}=0.5;"


"A^{-1} =\cfrac{1}{0.65}\cdot\begin{bmatrix} 0.5 & - 3 & 3\\0.05 & 1&-1\\-0.025&0.15&0.5\end{bmatrix};"


"\begin{bmatrix}
 x\\
 y\\
z
\end{bmatrix}=\cfrac{1}{0.65}\cdot\begin{bmatrix} 0.5 & - 3 & 3\\0.05 & 1&-1\\-0.025&0.15&0.5\end{bmatrix}\cdot\begin{bmatrix}
 4\\
 1\\
3
\end{bmatrix}="

"=\cfrac{1} {0.65} \cdot\begin{bmatrix}
 0.5\cdot4-3\cdot1+3\cdot3\\
 0.05\cdot4+1\cdot1-1\cdot3\\
-0.025\cdot4+0.15\cdot1+0.5\cdot3
\end{bmatrix}="

"=\cfrac{1} {0.65} \cdot\begin{bmatrix}
 8\\
 - 1.8\\
1.55
\end{bmatrix}=\begin{bmatrix}
 12.31\\
 - 2.77\\
2.38
\end{bmatrix}."